# Quantum many-body analysis of a spin-2 bosons with two-body inelastic decay

### v1

@@ Abstract
Bose-Einstein condensates (BECs) of $^{87}\textrm{Rb}$ atoms with a hyperfine spin of 2 are open quantum systems, where the atoms are lost through two-body inelastic collisions. In this dissipation process, a collision channel with total spin of 4 is forbidden by angular momentum conservation, which results in magnetization of the atoms remaining in the condensate. Here, we investigate the quantum many-body properties of spin-2 bosons that undergo two-body atomic loss. We show that the system finally reaches a steady state, which is a mixture of the states with maximum total spins. In addition, we find that a non-classical steady state can be obtained by quenching the quadratic Zeeman coefficient.
@@ I Introduction
Spin-2 BECs have attracted much interest because they exhibit more diverse magnetic phases than spin-1 BECs [13, 14, 15]. However, for the atomic species used in experiments (e.g., $^{87}\textrm{Rb}$), the spin-2 hyperfine state has a larger energy than the spin-1 state and is unstable against transition to the spin-1 state through binary collisions [16, 17, 18]. In such inelastic collisions, the hyperfine energy (e.g., $\simeq 6.8$ GHz for $^{87}\textrm{Rb}$) transforms to kinetic energy and the colliding atoms escape from the trap. Such a system with particle dissipation is an example of an open quantum system [19, 20, 21].
@@
In the experiment reported in Ref. [21], magnetization was observed to dynamically develop in a spin-2 BEC of $^{87}\textrm{Rb}$ atoms, even though the spin-dependent contact interaction was not ferromagnetic. The mechanism involved in such magnetization is spin-dependent atomic loss, in which two-body inelastic collisions of particles with a total spin of 4 (=2+2=2+2) is prohibited because of angular momentum conservation (i.e., only atomic pairs having different spin directions are lost). Consequently, the remaining atoms tend to become magnetized in the same direction.
@@
In Ref. [21], the experimental results have been confirmed by numerical calculations using the mean-field approximation. However, the mean-field picture is inappropriate for systems with a small number of particles. In fact, when the trap confinement is tight and all atoms occupy the same spatial wave function, the true ground state of a spin-1 BEC with antiferromagnetic interaction is a symmetry-preserving state referred to as a fragmented state [24, 22, 23]. Recently, such a quantum many-body state was experimentally observed in a BEC with $\sim 100$ atoms [25, 26].
In the present study, we perform quantum many-body analysis of spin-2 bosons that undergo two-body inelastic loss. We prepare an initial state in which all atoms occupy the $m=0$ magnetic sublevel, as in the experiment [21] and investigate the time evolution using the Lindblad master equation [27]. We show that the system eventually reaches a steady state, which is a statistical mixture of the states with maximum total spins (i.e., the final state is fully magnetized). We perform numerical simulations of the time evolution of the density matrix and investigate physical quantities in the final state. In the presence of the quadratic Zeeman effect, the transverse magnetization is found to be enhanced during the early stage of evolution of the system. When the quadratic Zeeman coefficient is quenched, the final steady state acquires a large transverse magnetization. This state exhibits a nonclassical magnetization distribution, indicating a Schrödinger-cat-like state.
@@
The remainder of this paper is organized as follows. First, we formulate the open quantum system in Sec. II and describe the dynamics of the quantum state using the Lindblad master equation. In Sec. III, we focus on the steady states of this equation and discuss their general properties. In Sec. IV, we present the results of a numerical analysis of the dynamics of the system. The conclusions to this study are provided in Sec. V.
II Formulation of problem
@@ II.1 Master equation with atomic loss
We consider spin-2 bosonic atoms confined in a trap potential $V(\bm{r})$. The field operators $\hat{\psi}_{m}(\bm{r})$ annihilate atoms in the magnetic sublevels m=2,1,0,1,2m=-2,-1,0,1,2, which satisfy the bosonic commutation relations $\hat{\psi}_{m}(\bm{r}),\hat{\psi}_{m^{\prime}}^{\dagger}(\bm{r}^{\prime})]=\delta(\bm{r}-\bm{r}^{\prime})\delta_{m,m^{\prime}}$. The non-intaracting component of the Hamiltonian is given by
@@
where $M$ is the mass of an atom and $q$ is the quadratic Zeeman coefficient. Here, we ignore the linear Zeeman effect because it can be eliminated by the unitary transformation.
The interaction term $\hat{H}_{\textrm{int}}$ that represents the elastic collisions is written as
$\hat{H}_{\textrm{int}}=\int\sum_{\mathcal{F}=0,2,4}\frac{g_{\mathcal{F}}}{2}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})d\bm{r},$ (2)
where $\mathcal{F}$ is the total spin of the colliding atoms and $g_{\mathcal{F}}=4\pi\hbar^{2}a_{\mathcal{F}}/M$ with $a_{\mathcal{F}}$ being the $s$-wave scattering length of the collision channel with $\mathcal{F}$. The operator $\mathcal{\hat{A}}_{\mathcal{FM}}$ is defined as
@@
where $C_{m,m^{\prime}}^{\mathcal{FM}}$ is the Clebsch-Gordan coefficient. The Hamiltonian for the total system is given by $\hat{H}=\hat{H}_{0}+\hat{H}_{\textrm{int}}$.
Next, we consider two-body inelastic collisions, in which a transition from a hyperfine spin-2 state to a hyperfine spin-1 state occurs. The energy difference between hyperfine spin-2 and spin-1 states for $^{87}\textrm{Rb}$ is $\simeq 6.8$ GHz, and this energy is transferred to the kinetic energy of the colliding atoms, causing them to escape from the trap. The dynamics of such an open quantum system with atomic loss can be described by the density matrix $\hat{\rho}$ as
$\frac{\partial\hat{\rho}}{\partial t}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}]+\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}},$ (4)
where $(\partial\hat{\rho}/\partial t)_{\textrm{loss}}$ represents the atomic loss due to inelastic collisions. Assuming that the atoms escaping from the trap do not interact with other atoms, the atomic loss is described by a Markov process. In this case, $(\partial\hat{\rho}/\partial t)_{\textrm{loss}}$ in Eq. (4) is given in the Lindblad form as [20]
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}}=\int\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left[2\hat{A}_{\mathcal{FM}}(\bm{r})\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\right$. (5)
$\left.-\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})\hat{\rho}-\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})\right]d\bm{r},$
where $b_{\mathcal{F}}$ is the loss coefficient for a collision channel with total spin $\mathcal{F}$. Notably, for $\mathcal{F}=4$, inelastic collisions are forbidden (i.e., $b_{4}=0$) because of the conservation of spin angular momentum. The first term in the square bracket in Eq. (5) is called a jump term, which eliminates two particles from the system.
@@ II.2 Previous results
Here, we briefly review the results of a previous study on a spin-2 $^{87}{\rm Rb}$ BEC with particle dissipation [21]. In this study, all atoms were initially in the $m=0$ state and the subsequent dynamics was investigated. It was observed that the system eventually became fully magnetized in the transverse direction despite the ferromagnetic state being energetically unfavorable [15]. This is due to the spin-dependent inelastic collisions ($b_{4}=0$); that is, atoms with different spin directions are selectively lost.
A numerical analysis was also performed in Ref. [21], in which Eq. (4) was reduced to the dissipative Gross-Pitaevskii equation. In the mean-field approximation, the macroscopic wave function $\psi_{m}(\bm{r})\equiv\textrm{Tr}[\hat{\psi}_{m}(\bm{r})\hat{\rho}$] obeys [20]
$i\hbar\frac{\partial\psi_{m}}{\partial t}=\left[-\frac{\hbar^{2}\nabla^{2}}{2M}+V(\bm{r})+qm^{2}\right]\psi_{m}(\bm{r})$ (6)
$+\sum_{\begin{subarray}{c}\mathcal{F}=0,2,4\\ \mathcal{M}=-\mathcal{F},...,\mathcal{F}\end{subarray}}\tilde{g}_{\mathcal{F}}\sum_{\begin{subarray}{c}i,j,k=-2\end{subarray}}^{2}C_{m,i}^{\mathcal{FM}}C_{j,k}^{\mathcal{FM}}\psi_{i}^{\star}(\bm{r})\psi_{j}(\bm{r})\psi_{k}(\bm{r}),$
@@ II.3 Single-mode approximation
We consider a situation in which the trap is tight and the size of the system is much smaller than the spin healing length. In this case, all particles occupy the same spatial wave function with a uniform spin state and we can use the single-mode approximation. The field operators can be rewritten as
$\hat{\psi}_{m}(\bm{r})\simeq\hat{a}_{m}\psi_{\textrm{SMA}}(\bm{r}),$ (7)
where $\hat{a}_{m}$ are bosonic operators and the wave function $\psi_{\textrm{SMA}}(\bm{r})$ is the ground state of $\hbar^{2}\nabla^{2}/(2M)+V]\psi_{\rm SMA}=\varepsilon_{0}\psi_{\rm SMA}$ satisfying the normalization condition $\int|\psi_{\textrm{SMA}}(\bm{r})|^{2}d\bm{r}=1$. Substituting Eq. (7) into Eqs. (1) and (2), we obtain
@@
$+\sum_{\mathcal{F}=0,2,4}\frac{g_{\mathcal{F}}}{2V^{\textrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}\hat{A}_{\mathcal{FM}}$,
where $\hat{N}=\sum_{m}\hat{a}_{m}^{\dagger}\hat{a}_{m}$, $V^{\rm eff}=\left[\int|\psi_{\textrm{SMA}}(\bm{r}))|^{4}d\bm{r}\right]^{-1}$, and
@@
Similarly, substituting Eq. (7) into Eq. (5) gives
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}}=\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4V^{\textrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left(2\hat{A}_{\mathcal{FM}}\hat{\rho}\hat{A}_{\mathcal{FM}}^{\dagger}\right$. (10)
[…]
III GENERAL FORMS OF STEADY STATES
This section provides a general discussion of the steady state for the Lindblad equation (4) with a single-mode approximation. The steady states satisfy $\partial\hat{\rho}/\partial t=0$. In the present case, the unitary and dissipation terms on the right-hand side of Eq. (4) must each be zero in a steady state.
First, let us consider the condition for
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{unitary}}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}]=0.$ (11)
Denoting the energy eigenstates of $\hat{H}$ by $|n\rangle$ and their corresponding energy eigenvalues by $E_{n}$, we can generally write the density matrix as
$\hat{\rho}=\sum_{n,m}c_{nm}|n\rangle\langle m|,$ (12)
where $c_{nm}$ is a positive semidefinite Hermitian matrix with unit trace. Equation (11) can be written as
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{unitary}}$ $=\frac{1}{i\hbar}\sum_{n,m}(E_{n}-E_{m})c_{nm}|n\rangle\langle m|$ (13)
$=0.$
Therefore, the density matrix that satisfies Eq. (11) is given by Eq. (12) with $E_{n}=E_{m}$ for all $n$ and $m$.
The energy eigenstates of $\hat{H}$ are expressed in the form $|N_{0},N_{S},F,F_{z}\rangle$ for $q=0$, where $N_{S}$ is the number of particles forming singlet pairs, $F$ is the total spin, and $F_{z}$ is the magnetic quantum number. The total number of particles is $N=N_{0}+2N_{S}$. The energy eigenvalue is given by [22]
$E=$ $\varepsilon_{0}N+\frac{c_{0}}{2}N(N-1)$ (14)
$+$ $\frac{c_{1}}{2}[F(F+1)-6N]+\frac{c_{2}}{5}N_{S}(N+N_{0}+3),$
@@ III GENERAL FORMS OF STEADY STATES
We note that the energy eigenstates $|N_{0},N_{S},F,F_{z}\rangle$ are independent of the values of $g_{0},g_{2}$ and $g_{4}$.
Second, let us consider the condition for
$\left(\frac{\partial\hat{\rho}}{\partial t}\right)_{\textrm{loss}}=0,$ (16)
where the left-hand side is given by Eq. (10). We will show that such states are mixtures of energy eigenstates with the maximum total spin $F$,

### v2

# Quantum many-body analysis of spin-2 bosons with two-body inelastic decay

@@ Abstract
Bose-Einstein condensates of $^{87}\mathrm{Rb}$ atoms with a hyperfine spin of 2 are open quantum systems, where the atoms are lost through two-body inelastic collisions. In this dissipation process, a collision channel with total spin of 4 is forbidden by angular momentum conservation, which results in magnetization of the atoms remaining in the condensate. Here, we investigate the quantum many-body properties of spin-2 bosons that undergo two-body atomic loss. We show that the system finally reaches a steady state, which is a mixture of the states with maximum total spins. In addition, we find that a nonclassical steady state can be obtained by applying and quenching the quadratic Zeeman field.
@@ I Introduction
Spin-2 BECs have attracted much interest because they exhibit more diverse magnetic phases than spin-1 BECs [13, 14, 15]. However, for the atomic species used in experiments (e.g., $^{87}\mathrm{Rb}$), the spin-2 hyperfine state has a larger energy than the spin-1 state and is unstable against transition to the spin-1 state through binary collisions [16, 17, 18]. In such inelastic collisions, the hyperfine energy (e.g., $\simeq 6.8$ GHz for $^{87}\mathrm{Rb}$) transforms to kinetic energy and the colliding atoms escape from the trap. Such a system with particle dissipation is an example of an open quantum system [19, 20, 21].
@@
In the experiment reported by Eto et al. [21], magnetization was observed to dynamically develop in a spin-2 BEC of $^{87}\mathrm{Rb}$ atoms, even though the spin-dependent contact interaction was not ferromagnetic. The mechanism involved in such magnetization is spin-dependent atomic loss, in which two-body inelastic collisions of particles with a total spin of 4 (=2+2=2+2) is prohibited because of angular momentum conservation (i.e., only atomic pairs with different spin directions are lost). Consequently, the remaining atoms tend to become magnetized in the same direction.
@@
In Ref. [21], the experimental results were confirmed by numerical calculations using the mean-field approximation. However, the mean-field picture is inappropriate for systems with a small number of particles. In fact, when the trap confinement is tight and all atoms occupy the same spatial wave function, the true ground state of a spin-1 BEC with antiferromagnetic interaction is a symmetry-preserving state referred to as a fragmented state [24, 22, 23]. Recently, such a quantum many-body state was experimentally observed in a BEC with $\sim 100$ atoms [25, 26].
In the present study, using the Lindblad master equation [27], we perform a quantum many-body analysis of spin-2 bosons that undergo two-body inelastic loss. We obtain two main results. The first result is that the steady state reached after longtime evolution is shown to be a statistical mixture of the states with maximum total spins (i.e., the final state is fully magnetized). The second result is that we can obtain a nonclassical quantum many-body state (Schrödinger-cat-like state) as the steady state. By applying a quadratic Zeeman field and quenching it, we can increase the probability of obtaining the nonclassical state. We numerically solve the time evolution of the master equation for $20$ particles and demonstrate the two aforementioned phenomena.
@@
The remainder of this paper is organized as follows. First, we formulate the present open quantum system in Sec. II and describe the dynamics of the system using the Lindblad master equation. In Sec. III, we focus on the steady states of this equation and analyze their general properties. In Sec. IV, we present the results of numerical simulations of the dynamics of the system. The conclusions to this study are provided in Sec. V.
II Formulation of the problem
@@ II.1 Master equation with atomic loss
We consider spin-2 bosonic atoms confined in a trap potential $V(\bm{r})$. The field operators $\hat{\psi}_{m}(\bm{r})$ annihilate atoms in the magnetic sublevels m=2,1,0,1,2m=-2,-1,0,1,2, which satisfy the bosonic commutation relations $\hat{\psi}_{m}(\bm{r}),\hat{\psi}_{m^{\prime}}^{\dagger}(\bm{r}^{\prime})]=\delta(\bm{r}-\bm{r}^{\prime})\delta_{m,m^{\prime}}$. The non-interacting component of the Hamiltonian is given by
@@
where $M$ is the mass of an atom and $q$ is the quadratic Zeeman coefficient. Here, we ignore the linear Zeeman effect because it can be eliminated by the unitary transformation. The quadratic Zeeman term is realized by applying an external magnetic field. The interaction term $\hat{H}_{\mathrm{int}}$ that represents the elastic collisions is written as
$\hat{H}_{\mathrm{int}}=\int\sum_{\mathcal{F}=0,2,4}\frac{g_{\mathcal{F}}}{2}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})d\bm{r},$ (2)
where $\mathcal{F}$ is the total spin of the colliding atoms and $g_{\mathcal{F}}=4\pi\hbar^{2}a_{\mathcal{F}}/M$ with $a_{\mathcal{F}}$ being the $s$-wave scattering length of the collision channel with $\mathcal{F}$. The operator $\hat{A}_{\mathcal{FM}}$ is defined as
@@
where $C_{m,m^{\prime}}^{\mathcal{FM}}$ is the Clebsch-Gordan coefficient. The Hamiltonian for the total system is given by $\hat{H}=\hat{H}_{0}+\hat{H}_{\mathrm{int}}$.
Next, we consider two-body inelastic collisions, in which a transition from a hyperfine spin-2 state to a hyperfine spin-1 state occurs. The energy difference between hyperfine spin-2 and spin-1 states for $^{87}\mathrm{Rb}$ is $\simeq 6.8$ GHz, and this energy is transferred to the kinetic energy of the colliding atoms, causing them to escape from the trap. The dynamics of such an open quantum system with atomic loss can be described by the density operator $\hat{\rho}(t)$ as
$\frac{d\hat{\rho}(t)}{dt}=\frac{1}{i\hbar}[\hat{H},\hat{\rho}(t)]+\left(\frac{d\hat{\rho}(t)}{dt}\right)_{\mathrm{loss}},$ (4)
where $(d\hat{\rho}(t)/dt)_{\mathrm{loss}}$ represents the atomic loss due to inelastic collisions. Assuming that the atoms escaping from the trap do not interact with other atoms, the atomic loss is described by a Markov process. In this case, $(d\hat{\rho}(t)/dt)_{\mathrm{loss}}$ in Eq. (4) is given in the Lindblad form as [20]
$\left(\frac{d\hat{\rho}(t)}{dt}\right)_{\mathrm{loss}}=\int\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left[2\hat{A}_{\mathcal{FM}}(\bm{r})\hat{\rho}(t)\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\right$. (5)
$\left.-\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})\hat{\rho}(t)-\hat{\rho}(t)\hat{A}_{\mathcal{FM}}^{\dagger}(\bm{r})\hat{A}_{\mathcal{FM}}(\bm{r})\right]d\bm{r},$
where $b_{\mathcal{F}}$ is the loss coefficient for a collision channel with total spin $\mathcal{F}$. Notably, for $\mathcal{F}=4$, inelastic collisions are forbidden (i.e., $b_{4}=0$) because of the conservation of spin angular momentum. The first term in the square brackets in Eq. (5) is called a jump term, which eliminates two particles from the system.
@@ II.2 Previous results
Here, we briefly review the results of a previous study on a spin-2 $^{87}\mathrm{Rb}$ BEC with particle dissipation [21]. In this previous study, all atoms were initially in the $m=0$ state and the subsequent dynamics was investigated. It was observed that the system eventually became fully magnetized in the transverse direction despite the ferromagnetic state being energetically unfavorable [15]. This is due to the spin-dependent inelastic collisions ($b_{4}=0$); that is, atoms with different spin directions are selectively lost.
A numerical analysis was also performed in Ref. [21], in which Eq. (4) was reduced to the dissipative Gross-Pitaevskii equation. In the mean-field approximation, the macroscopic wave function $\psi_{m}(\bm{r},t)\equiv\mathrm{Tr}[\hat{\psi}_{m}(\bm{r})\hat{\rho}(t)]$ obeys [20]
$i\hbar\frac{\partial\psi_{m}(\bm{r},t)}{\partial t}=\left[-\frac{\hbar^{2}\nabla^{2}}{2M}+V(\bm{r})+qm^{2}\right]\psi_{m}(\bm{r},t)$ (6)
$+\sum_{\begin{subarray}{c}\mathcal{F}=0,2,4\\ \mathcal{M}=-\mathcal{F},\dots,\mathcal{F}\end{subarray}}\tilde{g}_{\mathcal{F}}\sum_{\begin{subarray}{c}i,j,k=-2\end{subarray}}^{2}C_{m,i}^{\mathcal{FM}}C_{j,k}^{\mathcal{FM}}\psi_{i}^{\star}(\bm{r},t)\psi_{j}(\bm{r},t)\psi_{k}(\bm{r},t),$
@@ II.3 Single-mode approximation
We consider a situation in which atoms are tightly trapped and the size of the atomic cloud is much smaller than the spin healing length. In this case, all atoms occupy the same spatial wave function with a uniform spin state and we can use the single-mode approximation. The field operators can be rewritten as
$\hat{\psi}_{m}(\bm{r})\simeq\hat{a}_{m}\psi_{\mathrm{SMA}}(\bm{r}),$ (7)
where $\hat{a}_{m}$ are bosonic operators and the wave function $\psi_{\mathrm{SMA}}(\bm{r})$ is the ground state of $\hbar^{2}\nabla^{2}/(2M)+V]\psi_{\rm SMA}=\varepsilon_{0}\psi_{\rm SMA}$ satisfying the normalization condition $\int|\psi_{\mathrm{SMA}}(\bm{r})|^{2}d\bm{r}=1$. Substituting Eq. (7) into Eqs. (1) and (2), we obtain
@@
$+\sum_{\mathcal{F}=0,2,4}\frac{g_{\mathcal{F}}}{2V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\hat{A}^{\dagger}_{\mathcal{FM}}\hat{A}_{\mathcal{FM}}$,
where $\hat{N}=\sum_{m}\hat{a}_{m}^{\dagger}\hat{a}_{m}$, $V^{\rm eff}=\left[\int|\psi_{\mathrm{SMA}}(\bm{r})|^{4}d\bm{r}\right]^{-1}$, and
@@
Similarly, substituting Eq. (7) into Eq. (5) gives the single-mode approximation of the particle loss terms. We thus obtain the master equation for the single-mode approximation as
$\frac{d\hat{\rho}(t)}{dt}$ $=\frac{1}{i\hbar}[\hat{H},\hat{\rho}(t)]$ (10)
$+\sum_{\mathcal{F}=0,2}\frac{b_{\mathcal{F}}}{4V^{\mathrm{eff}}}\sum_{\mathcal{M}=-\mathcal{F}}^{\mathcal{F}}\left[2\hat{A}_{\mathcal{FM}}\hat{\rho}(t)\hat{A}_{\mathcal{FM}}^{\dagger}\right$.
$\left.-\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}(t)-\hat{\rho}(t)\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\right$],
where $\hat{H}$ is given by Eq. (8).
The energy eigenstates of $\hat{H}$ in Eq. (8) are expressed in the form $|N_{0},N_{S},F,F_{z}\rangle$ for $q=0$, where $N_{S}$ is the number of particles forming singlet pairs, $N_{0}$ is the number of remaining particles, $F$ is the total spin, and $F_{z}$ is the magnetic quantum number [22]. The total number of particles is $N=N_{0}+2N_{S}$. The energy eigenvalue is given by
$E=$ $\varepsilon_{0}N+\frac{c_{0}}{2V^{\mathrm{eff}}}N(N-1)$ (11)
$+$ $\frac{c_{1}}{2V^{\mathrm{eff}}}[F(F+1)-6N]+\frac{c_{2}}{5V^{\mathrm{eff}}}N_{S}(N+N_{0}+3),$
@@
We note that the energy eigenstates $|N_{0},N_{S},F,F_{z}\rangle$ are independent of the values of the interaction parameters $g_{0}$, $g_{2}$, and $g_{4}$.
III STEADY STATES REACHED AFTER LONG TIME
This section provides a general discussion of the steady state for the master equation (10) with a single-mode approximation. The steady states satisfy
$\frac{d\hat{\rho}(t)}{dt}=0.$ (13)
In Sec. III.1, we investigate the form of the steady state for a system with an initial particle number of two; in Sec. III.2, we extend the discussion to general particle numbers.
III.1 Two-particle solution
As a simple toy model, we consider an initial state in which two atoms occupy the $m=0$ state:
$|0,0,2,0,0\rangle=\frac{1}{\sqrt{2}}(\hat{a}_{0}^{\dagger})^{2}|\mathrm{vac}\rangle,$ (14)
where $|\mathrm{vac}\rangle$ is the vacuum state. We solve the master equation (10) and obtain the steady state for $t\to\infty$.
We abbreviate the relevant energy eigenstates as
$|0\rangle$ $\equiv|\mathrm{vac}\rangle,$ (15)
$|1\rangle$ $\equiv|N_{0}=0,N_{S}=1,F=0,F_{z}=0\rangle,$
$|2\rangle$ $\equiv|N_{0}=2,N_{S}=0,F=2,F_{z}=0\rangle,$
$|3\rangle$ $\equiv|N_{0}=2,N_{S}=0,F=4,F_{z}=0\rangle.$
The initial state in Eq. (14) can be expanded by these energy eigenstates as
$|0,0,2,0,0\rangle=\frac{1}{\sqrt{35}}\left(\sqrt{7}|1\rangle-\sqrt{10}|2\rangle+3\sqrt{2}|3\rangle\right).$ (16)
We write the time evolution of the density operator as
$\hat{\rho}(t)=\sum_{i,j=0}^{3}c_{ij}(t)|i\rangle\langle j|.$ (17)
Applying $\hat{A}_{\mathcal{FM}}$ to the states $|i\rangle$, we have $\hat{A}_{00}|1\rangle=\hat{A}_{20}|2\rangle=\sqrt{2}|0\rangle$ with all other combinations being zero; their Hermitian conjugates give $\hat{A}_{00}^{\dagger}|0\rangle=\sqrt{2}|1\rangle$ and $\hat{A}_{20}^{\dagger}|0\rangle=\sqrt{2}|2\rangle$. Consequently, Eq. (10) is expressed in the form
$\frac{d\hat{\rho}(t)}{dt}=$ $\frac{1}{i\hbar}\sum_{i,j=0}^{3}\left[E_{i}-E_{j}-i(D_{i}+D_{j})\right]c_{ij}(t)|i\rangle\langle j|$ (18)
$+2D_{1}c_{11}(t)|0\rangle\langle 0|+2D_{2}c_{22}(t)|0\rangle\langle 0|,$
where $E_{i}$ and $D_{i}$ are $E_{0}=D_{0}=0,E_{1}=g_{0}/V^{\mathrm{eff}},D_{1}=b_{0}/(2V^{\mathrm{eff}}),E_{2}=g_{2}/V^{\mathrm{eff}},D_{2}=b_{2}/(2V^{\mathrm{eff}})$, $E_{3}=g_{4}/V^{\mathrm{eff}}$, and $D_{3}=0$.
We solve Eq. (18) for the initial condition of $\hat{\rho}(0)=|0,0,2,0,0\rangle\langle 0,0,2,0,0|$. For $i\neq j$, we have
$\frac{dc_{ij}(t)}{dt}=\frac{1}{i\hbar}[E_{i}-E_{j}-i(D_{i}+D_{j})]c_{ij}(t),$ (19)
giving $c_{ij}(t)=c_{ij}(0)\exp\{[E_{i}-E_{j}-i(D_{i}+D_{j})]t/(i\hbar)\}$. Since $D_{i}+D_{j}>0$ for $i\neq j$, the off-diagonal coefficients become $c_{ij}(t)\to 0$ as $t\to\infty$. For the diagonal elements $c_{ii}$, Eq. (18) gives
$\frac{dc_{00}(t)}{dt}$ $=2D_{1}c_{11}(t)+2D_{2}c_{22}(t),$ (20)
$\frac{dc_{11}(t)}{dt}$ $=-2D_{1}c_{11}(t),$
$\frac{dc_{22}(t)}{dt}$ $=-2D_{2}c_{22}(t),$
$\frac{dc_{33}(t)}{dt}$ $=0,$
which are solved to give
$c_{00}(t)$ $=c_{11}(0)(1-e^{-2D_{1}t})+c_{22}(0)(1-e^{-2D_{2}t}),$ (21)
$c_{11}(t)$ $=c_{11}(0)e^{-2D_{1}t},$
$c_{22}(t)$ $=c_{22}(0)e^{-2D_{2}t},$
$c_{33}(t)$ $=c_{33}(0).$
In the limit of $t\to\infty$, $c_{11}(t)\rightarrow 0$ and $c_{22}(t)\rightarrow 0$, whereas $c_{33}(t)$ maintains its initial value. Thus, we obtain the steady state as
$\hat{\rho}(t)\rightarrow\frac{17}{35}|0\rangle\langle 0|+\frac{18}{35}|3\rangle\langle 3|~~~~(t\rightarrow\infty).$ (22)
Other than the vacuum state $|0\rangle$, this steady state only contains the state $|3\rangle=|N_{0}=2,N_{S}=0,F=4,F_{z}=0\rangle$ that has the maximum total spin $F=2N=4$. This is because particle loss occurs only through the ${\cal F}=0$ and ${\cal F}=2$ channels. The maximally polarized state therefore does not undergo particle loss and survives in the steady state. The above simple analysis for two particles implies that for general numbers of particles also, the steady state only contains the states with maximum total spins. In the next subsection, we will show that this is the case.
III.2 General forms of steady states
In this subsection, we prove that the steady state satisfying Eq. (13) is a mixture of energy eigenstates with the maximum total spin. In the following, we denote the density operator for the $N$-particle subspace as $\hat{\rho}_{N}$ and express the density operator for the total Hilbert space as
$\hat{\rho}(t)=\oplus_{N=0}^{N_{\mathrm{ini}}}\hat{\rho}_{N}(t),$ (23)
where $N_{\mathrm{ini}}$ is the maximum number of atoms contained in the initial state.
First, we consider the case of $q=0$. We abbreviate the states with the maximum total spin $F=2N$ as